OpenStudy (anonymous):

Can someone help me please... y+2=-(x-1) change to slope intercept

OpenStudy (anonymous):

change to slope intercept means change to the form y=mx+c first expand the bracket and change to move all to the variables to the right whilst making y the subject what do you get?

OpenStudy (fibonaccichick666):

^That means, first use the distributive property on the right side to simplify \(-(x-1)\).

OpenStudy (anonymous):

@FibonacciChick666 so x becomes negative and 1 becomes positive???????? and it would be y+2=-x+1

OpenStudy (fibonaccichick666):

that is a correct first step.

OpenStudy (fibonaccichick666):

good job, so now, you want to get that y by itself, how can you "undo" the addition of a 2?

OpenStudy (anonymous):

@FibonacciChick666 so I would subtract 2 from both sides??

OpenStudy (fibonaccichick666):

that is correct! show me what you get when you do

OpenStudy (anonymous):

@FibonacciChick666 I got Y=-x-1 I not sure if its correct though.

OpenStudy (fibonaccichick666):

that is correct!

OpenStudy (anonymous):

@FibonacciChick666 Yay I did it.. haha thank you :)

OpenStudy (fibonaccichick666):

np! Thank you for working through it!